OpenStudy (anonymous):

HELP? FIND ALL SOLUTIONS OF THE EQUATION 4SIN^2X=4COS x+5 IN THE INTERVAL [0,2PI)

OpenStudy (callisto):

\[4sin^2x=4cosx+5\]\[4(1-cos^2x)=4cosx+5\]\[4-4cos^2x=4cosx+5\]\[4cos^2x + 4cosx+1=0\]\[(2cosx+1)^2=0\]\[cosx=-0.5\]Can you solve x?

OpenStudy (anonymous):

okay can u please explain step-by-step please? i know its a drag but can u?

OpenStudy (callisto):

It's okay :) First, you can see that there is a \(sin^2x\) on the left and a cosx on the right. To solve the equation, we'd better change it into one ratio. cosx is difficult to change. So, I changed \(sin^2x\) into \(1-cos^2x\) by using the formula \(sin^2x + cos^2x =1\) Got it so far?

OpenStudy (anonymous):

YEAH :)

OpenStudy (callisto):

Then, it becomes the quadratic equation... Any questions?

OpenStudy (anonymous):

well yeah... why exactly does it become the quad. equation?

OpenStudy (anonymous):

like as the next step for this problem

OpenStudy (callisto):

because you can rearrange the terms into the form \(ax^2 + bx + c =0\) See the fourth line there~

OpenStudy (anonymous):

okay now im getting a little thrown off :(

OpenStudy (callisto):

Hmm... which part you don't understand?

OpenStudy (anonymous):

ummm welll they part where u say that< the terms could be rearranged into the form ax^2+bx+c=0> why is this equation necessary 4 this problem?

OpenStudy (anonymous):

or better yet at this point in the problem?

OpenStudy (callisto):

Because you need to solve it. When you see quadratic question, you can solve it easier..

OpenStudy (callisto):

Actually, it doesn't matter if it is a quadratic equation. As long as you can solve it, that's fine. But in this case, it just comes to be a quadratic equation ...

OpenStudy (anonymous):

okay so the quadartic equation comes out 2 b? what exactly.

OpenStudy (callisto):

See the second to the fourth lines :)

OpenStudy (anonymous):

okay so line four....which is 4cos^2+4cosx+1=0 this was gathered from subtracting the four and five and then multiplying it by -1 right and moving it over?

OpenStudy (callisto):

\[4-4cos^2x=4cosx+5\]\[4-4cos^2x-4=4cosx+5-4\]\[-4cos^2x=4cosx+1\]\[-4cos^2x+4cos^2x=4cosx+1+4cos^2x\]\[0=4cosx+1+4cos^2x\]\[4cos^2x+4cosx+1=0\]

OpenStudy (anonymous):

ok much clearer :] thank u!

OpenStudy (callisto):

welcome :) Any questions?

OpenStudy (anonymous):

yeah after this like there is nothing else? this is the answer for the problem?

OpenStudy (callisto):

Nope :| you need to solve it!!! Factor the left side, you'll get \[(2cosx+1)^2=0\]So, \[2cosx+1=0\]Solve x.

OpenStudy (anonymous):

so it'll be cosx=-1/2 and then i just find the two references on the unit circle right?

OpenStudy (callisto):

Yes Yes!

OpenStudy (anonymous):

k thank u!

OpenStudy (callisto):

welcome :)